OpenStudy (anonymous):

For the function f(x), write a function g(x) for the transformations ~~ f(x)=x; right 2 units

OpenStudy (anonymous):

To translate a function right n units, we replace \(x\) with \(x - n\). In this case we replace \(x\) with \(x-2\) \[ g(x) = f(x-2) = x-2 \]

OpenStudy (anonymous):

Ok, i understand now! Thank you! Does the same thing happen when moving left 2 as well???

OpenStudy (anonymous):

When moving right \(n\) units, you replace \(f(x)\) with \(f(x−n)\) When moving left \(n\) units, you replace \(f(x)\) with \(f(x+n)\) When moving down \(n\) units, you replace \(f(x)\) with \(f(x)-n\) When moving up \(n\) units, you replace \(f(x)\) with \(f(x)+n\) These are the steps for all translation (shifting) transformations.

OpenStudy (anonymous):

Wow, this helped me so much! Thank you!